# Understanding LLMs’ Cross-Lingual Context Retrieval: How Good It Is And Where It Comes From

### v1

@@ Abstract
The ability of cross-lingual context retrieval is a fundamental aspect of cross-lingual alignment of large language models (LLMs), where the model extracts context information in one language based on requests in another language. Despite its importance in real-life applications, this ability has not been adequately investigated for state-of-the-art models. In this paper, we evaluate the cross-lingual context retrieval ability of over 40 LLMs across 12 languages to understand the source of this ability, using cross-lingual machine reading comprehension (xMRC) as a representative scenario. Our results show that several small, post-trained open LLMs show strong cross-lingual context retrieval ability, comparable to closed-source LLMs such as GPT-4o, and their estimated oracle performances greatly improve after post-training. Our interpretability analysis shows that the cross-lingual context retrieval process can be divided into two main phases: question encoding and answer retrieval, which are formed in pre-training and post-training, respectively. The phasing stability correlates with xMRC performance, and the xMRC bottleneck lies at the last model layers in the second phase, where the effect of post-training can be evidently observed. Our results also indicate that larger-scale pretraining cannot improve the xMRC performance. Instead, larger LLMs need further multilingual post-training to fully unlock their cross-lingual context retrieval potential.  [^1] Our code and is available at https://github.com/NJUNLP/Cross-Lingual-Context-Retrieval.
@@ 1 Introduction
Since the rise of Large language models (LLMs), many models have demonstrated their strong capability in various NLP tasks Chang et al. (2024), e.g. ChatGPT [^2] https://chatgpt.com/, Claude [^3] https://claude.ai/, Gemini Gemini Team et al. (2024), LLaMA Grattafiori et al. (2024), Qwen Qwen et al. (2025), DeepSeek DeepSeek-AI et al. (2024b), etc. However, due to the dominance of English training data, most of these LLMs show their best performance in English Lai et al. (2023b); Wang et al. (2023). To improve their performance and efficiency in non-English languages, cross-lingual alignment has become a major research topic for multilingual LLMs Qi et al. (2023); Gao et al. (2024), which encourages LLMs to share capabilities across languages. For example, given requests with the same semantics but in different languages, LLMs should give consistent answers.
One fundamental aspect of such cross-lingual alignment is cross-lingual context retrieval, where the model needs to extract context information in the source language, e.g. English, to answer requests in the target language. However, there is yet no comprehensive evaluation and exploration of this ability.
In this paper, we evaluate the cross-lingual context retrieval ability of SOTA multilingual LLMs, and analyze the source and bottleneck of such capability. We use cross-lingual machine reading comprehension (xMRC) Cui et al. (2019) as a simplified but representative scenario, where the target knowledge to be retrieved is literally included in the context, and the model only has to copy part of the context as the answer. This ablates the need for knowledge storage and multilingual text generation, thus better focusing the research on cross-lingual context retrieval. Our main findings are:
  • Several SOTA post-trained open-source LLMs, especially in their 7-9B versions show strong xMRC ability, catching up with closed-source models. Larger models show higher English MRC performance, but larger gap between English and non-English.
  • Post-training significantly improves the estimated oracle performance to a very high level in all tested languages, setting space for improvement of real performance, while the effect of extended pretraining is minor.
  • The xMRC process can be divided into two phases within the model: question encoding and answer retrieval. The former forms in pre-training, and its stability correlates with the base model capability, as well as xMRC performance; the latter forms in post-training, and serves as a bottleneck for xMRC.
@@ 2.1.1 Scenarios
We use English to non-English (en-x) as a typical and common cross-lingual scenario, where the context and answer are in English, and the question is in non-English. Figure 1(a) shows a testing sample of en-x task across 3 example languages.
Meanwhile, we use non-English monolingual (x-x) as a comparative to the en-x scenario, in order to distinguish xMRC ability with monolingual MRC ability. Figure 1(b) shows a testing sample of x-x task in the same 3 example languages.
@@ 2.1.2 Metrics
We use the F1 score and exact match (EM) to evaluate xMRC performance, as adopted by XQuAD Artetxe et al. (2020). Because the models tend to include unnecessary source text in their correct answers, EM will under-estimate the correctness of the models’ response. As a result, we use F1 as the main performance metric.
@@
To measure the level of the models’ performance alignment between English and non-English languages, we calculate the average performance on all the non-English languages (en-x, x-x), as well as the ratio of non-English performances over English ones (en-x / en-en, x-x / en-en).
@@ 2.1.3 Performance bottleneck analysis
We distinguish content error in the model predictions from other less important types of error. Based on preliminary observations, we choose several major types of distracting errors:
@@
  • Gibberish: giving irrelevant, non-sense or hallucination output;
  • Refusal: refusing to answer the question;
  • Blank: providing no answer at all.
The last three error types can be aggregated as generation failure errors. We measure the proportion of these errors in the models’ incorrect responses. For example, the language error rate for test setting en-x is calculated as:
@@
where $r$ is the reference answer, $a$ is the model prediction, and $\mathrm{Lang}(\cdot)$ is a language detector for given text. Meanwhile, a generation failure error rate, e.g. gibberish, will be:
@@
where $\mathrm{Type}(\cdot)$ is an LLM-based error type classifier for generation failure samples.
@@
Because the xMRC score can be affected by errors in the generation process, it may not reflect the full potential of the model’s cross-lingual context retrieval ability. To ablate this effect, we estimate the oracle retrieval performance of the models by perturbation-based attribution [^4] We use Captum to perform the attribution (https://github.com/pytorch/captum) on contextual sentences or spans, which calculates the importance of each of them to the output. If the sentence/span with the correct answer in it receives the largest importance, we consider the model’s oracle retrieval on this testing sample is correct. Then, we calculate the accuracy of such selection respectively with one generation step or the whole generation process as the attributed target.
2.2 Analysis methods
@@ 2.2.1 Layer-wise attribution to reflect forward process
To better understand the forward calculation process of models performing xMRC retrieval, we need layer-wise attribution in order to see from which input part the LLM draws information into the residual stream Voita et al. (2024) at each layer.
Previous studies have proposed multiple layer-wise attribution methods, including attention-based Hao et al. (2021); Ferrando et al. (2022), backpropagation-based Voita et al. (2021), and decomposition-based Yang et al. (2023a) etc. Here we choose AttentionLRP Achtibat et al. (2024), because it can attribute latent representations in each model layer and is especially suitable for transformer models.
Based on this, we define the Major Relevance Depth (MRD) to estimate the maximum depth to which a token representation $x$ needs to be encoded, by calculating the layer number corresponding to the 95th percentile of its attributed relevance to model $m$’s output:
@@
[…]
where $r_{\text{out}}(x,i)$ refers to the normalized relevance score of token representation $x$ in layer $i$ to the final output given by AttentionLRP, which is based on Taylor decomposition on the attention, FFN, and normalization functions. Note that $N$ is one less than the total layer number (e.g. 31 for LLaMA-3.1-8B) because we attribute the output of the last layer to its precedent layers. If the MRD of an input token is $\hat{n}$, then we estimate that the token’s information participates in the context retrieval process only in the first $\hat{n}$-th layers. Then, for parts of the input, i.e., task description, demonstrations, question and context, we take the maximum MRD of tokens in each part to represent them, and calculate the mean MRD for each part on a group of testing samples to reflect the general distribution.
@@ 2.2.2 Hidden state similarity to measure cross-lingual alignment
[…]
@@
where $\overline{\text{Sim}}$ denotes the mean cosine similarity, $E$ denotes the hidden states from English MRC tasks, and $X$ denotes the hidden states from en-x xMRC tasks. $K$ is the total number of testing samples, $e_{k}$ and $x_{k}$ are the hidden states from the $k$-th parallel sample pair between English and language x, $x_{i},x_{j}$ are the hidden states from every two different samples in language x, and the cosine similarity $\text{Sim}(x,y)=x\cdot y/|x||y|$.
Name Modes Sizes
LLaMA-2 Chat 7B
LLaMA-3.1 Base / Instruct / Tuned 8B / 70B
Mistral-V0.3 Base / Instruct 7B
Qwen-2.5 Base / Instruct 7B / 72B
Gemma-2 Base / IT 9B
DeepSeek-V2 Base / Chat 16B
DeepSeek-V3 Chat 671B (37B Active)
GPT-3.5-Turbo - -
GPT-4o - -
Table 1: Selected models for main evaluation and analysis.
@@
Table 2: 2-shot F1 scores on en-x and x-x tasks, and 2-shot language error and generation failure error rates (%) on en-x tasks.
@@ 3.1 Dataset
We use the XQuAD dataset Artetxe et al. (2020) to measure the xMRC performance of LLMs, because its testing samples are parallel in all the 12 included languages and thus suitable for cross-lingual transforming. The XQuAD dataset is composed of 1190 parallel data points for each of its 12 languages, ensuring a consistent evaluation setup across languages. With an average context length of 702.50, the dataset provides sufficiently rich contexts for machine reading comprehension. Furthermore, the 12 languages cover diverse language families, scripts, and resource levels, making our evaluation more trustworthy and representative.
@@ 3.2 Models and tools
We adopt a variety of SOTA open and business LLMs, including LLaMA-3.1 Grattafiori et al. (2024), Mistral Jiang et al. (2023), Qwen-2.5 Qwen et al. (2025), Gemma-2 Team et al. (2024), DeepSeek V2&3 DeepSeek-AI et al. (2024a, b), GPT-3.5, and GPT-4o, in smaller and larger sizes. Table 1 shows our selected model list in the main evaluation and analysis results. Also, Table 4 in Appendix A.1 shows a full list of all the tested models, where some of them are of older versions and alternative sizes compared with the main-list models.
To verify the effect of post-training, we tune the LLaMA-3.1-8B model using the TULU-v3 dataset Lambert et al. (2025), which is a multilingual general instruction dataset, into a model called LLaMA-3.1-Tuned-8B (Appendix B.3 for details).
We also adopt tools for error type identification. For language error detection, we use Lingua  [^5] https://github.com/pemistahl/lingua with its high-accuracy mode, the accuracy of which is satisfactory in our tested languages. For generation failure errors detection, we use Qwen-2.5-72B-Instruct (prompt shown in Appendix A.3) to act like a classifier, distinguishing normal outputs and the three types of error.
@@ 3.3 Prompts
In our xMRC evaluation, we try our best to adopt the most suitable prompt templates for each tested model to minimize its limitation of performance. We try two prompting formats for each model and take the one with higher xMRC performance to represent that model. As shown in Appendix A.3, the v1 prompt format places the task description before the demonstrations, and the v2 prompt format places the task description after the demonstrations.
We also test each model in 2-shot and 0-shot settings. Because most of the tested LLMs perform better in the 2-shot setting, we focus our evaluation and analysis mainly on this setting. The 0-shot evaluation results are in Appendix A.2.
@@ 4.1 Evaluation results
The left part of Table 2 summarizes the F1 scores of our main-list models on both en-x cross-lingual MRC (xMRC) and x-x monolingual MRC tasks. Comparing the xMRC scores with the monolingual scores highlights the difference in model performance between cross-lingual and monolingual settings. Detailed F1 scores for each model, language, and task are further illustrated in Table 7 in Appendix A.2.
@@ 4.1.1 Cross-lingual performance
Generally, one can see the English MRC performance of most tested models are high (over 70 out of 100), but the xMRC scores ranges (from 45 to 78), showing the performance gap in context retrieval with English and non-English queries, even with the same English context and answer.
Down into individual models, while GPT-4o shows the highest cross-lingual performance and smallest language gap, several open-sourced post-trained model series, such as Gemma-2-IT, Qwen-2.5-Instruct and LLaMA-3.1-Instruct, also show comparable performance levels and small language gaps, indicating that newest training techniques contribute to higher xMRC performance.
An interesting observation is that, for LLaMA-3.1-8B and Gemma-2-9B, the English performances after post-training are close to those of the base versions, but the cross-lingual performances greatly improve. Also, this phenomenon is more prominent in smaller (7-9B) than larger models, bring the former a smaller performance gap between English and non-English. This brings interesting questions on the post-training of larger models.
@@ 4.1.2 Comparison with Monolingual performance
In general, the performance gaps between English and non-English monolingual MRC are much smaller for most of the tested LLMs than xMRC, and the Qwen models even show higher non-English performance than English. This suggests that non-English language fluency alone is not the main cause of the ranging xMRC performance.
Also, for base models and post-trained larger models ($\sim$70B), the monolingual performances in non-English languages are always higher than cross-lingual performance. A possible explanation to this may be the cross-lingual task is more difficult and less frequent in training, and it requires cross-lingual understanding.
However, for post-trained, smaller models (7-9B), the pattern flips, where the monolingual performances become consistently lower than cross-lingual ones. This result suggests that these models become more capable of using their English ability to assist context retrieval with non-English queries, overcoming the difficulty and low-frequency of the cross-lingual task. Also, since we should expect larger LLMs having better instruction following and understanding in general, this further highlights that post-training better elicits the cross-lingual context retrieval ability on smaller LLMs.
@@ 4.2.1 Error type ablation
Language error.
The right part of Table 2 shows the average percentage of language errors among all languages. One can see an expected advantage of post-trained models against base models, because the former are better in following the cross-lingual task format. However, since the error rate is low for all the post-trained models, it cannot be viewed as a bottleneck for the xMRC task.
Generation failure errors.
The right part of Table 2 also shows the average summed percentages of gibberish, refusal, and blank error in samples with wrong model answers (detailed percentages in Table 8-10(c) in Appendix A.4). One can see the generation failure rates are minor for most models, regardless of size and post-training, marking it not the bottleneck of xMRC either.

### v2

@@ Abstract
Cross-lingual context retrieval (extracting contextual information in one language based on requests in another) is a fundamental aspect of cross-lingual alignment, but the performance and mechanism of it for large language models (LLMs) remains unclear. In this paper, we evaluate the cross-lingual context retrieval of over 40 LLMs across 12 languages, using cross-lingual machine reading comprehension (xMRC) as a representative scenario. Our results show that post-trained open LLMs show strong cross-lingual context retrieval ability, comparable to closed-source LLMs such as GPT-4o, and their estimated oracle performances greatly improve after post-training. Our mechanism analysis shows that the cross-lingual context retrieval process can be divided into two main phases: question encoding and answer retrieval, which are formed in pre-training and post-training respectively. The phasing stability correlates with xMRC performance, and the xMRC bottleneck lies at the last model layers in the second phase, where the effect of post-training can be evidently observed. Our results also indicate that larger-scale pretraining cannot improve the xMRC performance. Instead, larger LLMs need further multilingual post-training to fully unlock their cross-lingual context retrieval potential.  [^1] Our code and results are available at https://github.com/NJUNLP/Cross-Lingual-Context-Retrieval
@@ 1 Introduction
Since the rise of Large language models (LLMs), many models have demonstrated their strong capability in various NLP tasks Chang et al. (2024), e.g. ChatGPT [^2] https://chatgpt.com, Claude [^3] https://claude.ai, Gemini Gemini Team et al. (2024), LLaMA Grattafiori et al. (2024), Qwen Qwen et al. (2025), DeepSeek DeepSeek-AI et al. (2024b), etc. However, due to the dominance of English training data, most of these LLMs show their best performance in English Lai et al. (2023b). To improve their performance and efficiency in non-English languages, cross-lingual alignment has become a major research topic for multilingual LLMs Qi et al. (2023); Wang et al. (2024), which encourages LLMs to share capabilities across languages. For example, given requests with the same semantics but in different languages, LLMs should give consistent answers.
Such alignment has been shown challenging when the task requires recalling trained knowledge Gao et al. (2024); Hu et al. (2025). Thus, a follow-up question is that, when the knowledge is presented in the context in one language (e.g. English), could LLMs retrieve it when answering requests in the another language? Note that this is different from multilingual context retrieval (the context and the request are in the same language) and cross-lingual information retrieval (retrieving queried text from a database). However, evaluation and mechanism analysis of this ability in LLMs are not fully explored.
In this paper, we evaluate cross-lingual context retrieval of SOTA multilingual LLMs, and analyze the mechanism of it. We use cross-lingual machine reading comprehension (xMRC) Cui et al. (2019) as a simplified but representative scenario because: (1) It is a representative NLP task where models need to retrieve cross-lingual knowledge within the context; (2) The target knowledge to be retrieved is entirely within the context, so minimal extra knowledge recall is needed; (3) The model only has to copy part of the context as the answer, simplifying evaluation and mitigating other factors such as low-resource language generation errors.
Furthermore, we conduct in-depth analysis of the performance bottleneck, oracle performance, and mechanism of LLMs doing the xMRC task, using suitable tools and specially designed metrics. Specifically, we propose a hypothesis of two-phased xMRC process in LLMs, and verify our hypothesis with experimental evidence.
Our main findings are:
  • Post-trained open-source LLMs, especially 7-9B versions, show strong xMRC ability, catching up with closed-source models. Larger models show higher English MRC performance, but larger gap between English and non-English.
  • Post-training significantly improves the estimated oracle performance to almost saturate in all tested languages, setting space for improvement of real performance, while the effect of extended pretraining is minor.
  • The xMRC process can be divided into two phases within the model: question encoding and answer retrieval. The former forms in pre-training, and its stability correlates with the base model capability, as well as xMRC performance; the latter forms in post-training, and serves as a bottleneck.
@@ 2.1.1 Scenarios
We use English to non-English (en-x) as a typical and common cross-lingual scenario, where the context and answer are in English, and the question is in non-English, because it avoids the requirement of non-English generation fluency that is not related to cross-lingual context retrieval, and keeps the answers always in English which makes answer verification easier and fairer. Meanwhile, non-English monolingual (x-x) is used as comparison to ablate the effect of non-English understanding. Figure 1 shows examples of the testing scenarios.
@@ 2.1.2 Metrics
We use the F1 score to evaluate xMRC performance, as adopted by XQuAD Artetxe et al. (2020). Exact match (EM) is not used because it tends to under-estimate the performance due to unnecessary source text in the answers (see Appendix A.1).
@@
To measure performance alignment between English and non-English, we calculate the average performance on all the non-English languages divided by English (en-x/en-en, x-x/en-en).
@@ 2.1.3 Performance bottleneck analysis
Based on observation of model responses, we distinguish error types as:
@@
  • Generation: generation errors including gibberish, refusal and blank answers;
  • Content: meaningful but incorrect answers in the correct language;
For example, the language error rate for test setting en-x is calculated as:
@@
where $r$ is the reference answer, $a$ is the model prediction, and $\mathrm{Lang}(\cdot)$ is a language detector for given text. Meanwhile, a generation error rate, e.g. gibberish, will be:
@@
where $\mathrm{Type}(\cdot)$ is an LLM-based error type classifier for generation errors.
@@
The xMRC score can be affected by errors in the generation process, under-estimating LLMs’ cross-lingual context retrieval ability. To ablate this effect, we estimate the oracle retrieval performance of LLMs by perturbation-based attribution [^4] We use Captum to perform the attribution (https://github.com/pytorch/captum) on contextual sentences or spans. If the sentence/span with the correct answer receives the largest attribution score, we consider the model’s oracle retrieval on this testing sample correct.
2.2 Mechanism analysis methods
@@ 2.2.1 Layer-wise attribution to reflect forward process
To better understand the forward process of LLMs performing xMRC, we need layer-wise attribution to observe information flow from context parts into the residual stream Voita et al. (2024) at each layer. Here we adopt AttentionLRP Achtibat et al. (2024), an attribution method based on Taylor Decomposition on attention, FFN and normalization modules of LLMs, that can calculate the relevance of token representations in each layer to the output.
Based on the attribution results, we define the Major Relevance Depth (MRD) to estimate the maximum depth to which a token representation $x$ needs to be encoded, by calculating the layer number corresponding to the 95th percentile of its attributed relevance to model $m$’s output:
@@
$\text{s.t.}\sum_{i=1}^{n}r_{\text{out}}(x,i)$ $\geqslant 0.95\sum_{i=1}^{N}r_{\text{out}}(x,i)$
where $r_{\text{out}}(x,i)$ refers to the normalized relevance score of token representation $x$ in layer $i$ to the final output given by AttentionLRP. A token-MRD of $\hat{n}$ indicates that the token information participates in the context retrieval process only in the first $\hat{n}-$th layers. Then, for parts of the input, i.e., task description, demonstrations, question and context, we take the maximum token-MRD of each to represent them, and calculate the mean part-MRD.
@@ 2.2.2 Hidden state similarity to measure cross-lingual alignment
To observe the cross-lingual alignment of internal xMRC process, we collect the hidden states in all model layers and calculate their cross-lingual similarity, and define a cross-lingual similarity ratio $S(\text{en,x})$ between English and language x:
$S(\text{en,x})$ $=\overline{\text{Sim}}(E,X)/\overline{\text{Sim}}(X,X)$
@@
where $\overline{\text{Sim}}$ denotes the mean cosine similarity, $\text{Sim}(x,y)=x\cdot y/|x||y|$. $E$ and $X$ denotes en-en and en-x hidden states. $K$ is the total number of samples, $e_{k}$ and $x_{k}$ are the hidden states from the $k$-th parallel sample pair between English and language x, $x_{i},x_{j}$ are the hidden states from every two different samples in language x.
@@
Table 1: 2-shot F1 scores on en-x and x-x tasks, and 2-shot language error and generation error rates (%) on en-x tasks.
@@ 3.1 Dataset
We use the XQuAD dataset Artetxe et al. (2020) to measure the xMRC performance of LLMs, because its testing samples are parallel in all the 12 included languages555en, de, es, vi, zh, hi, ar, el, ro, ru, th, tr and thus suitable for cross-lingual transforming. The dataset has 1190 parallel samples for each language and an average context length of 702.50 words. It also covers diverse language families, scripts, and resource levels.
@@ 3.2 Models and tools
We adopt a variety of SOTA open and business LLMs, including LLaMA-3.1 Grattafiori et al. (2024), Mistral Jiang et al. (2023), Qwen-2.5 Qwen et al. (2025), Gemma-2 Team et al. (2024), DeepSeek V2&3 DeepSeek-AI et al. (2024a, b), GPT-3.5, and GPT-4o, in smaller and larger sizes. Table 4 in Appendix B.1 shows a full list of all the tested models. We also tune the LLaMA-3.1-8B model with the TULU-v3 dataset Lambert et al. (2025) into a model called LLaMA-3.1-Tuned-8B (Appendix C.3 for details) to verify the effect of post-training.
For language error detection, we use Lingua [^6] https://github.com/pemistahl/lingua with its high-accuracy mode, the accuracy of which is satisfactory in our tested languages (see Appendix A.2). For generation errors detection, we use Qwen-2.5-72B-Instruct (prompt shown in Appendix A.3) to identify generation errors. To rule out the potential bias induced by Qwen judging itself, we adopt Gemini-2.5-Flash-Lite  [^7] https://gemini.google.com as another judge for cross-validation (see Appendix B.4).
@@ 3.3 Prompts
In our xMRC evaluation, we try two different prompt templates and use the one with higher performance for each model (see Appendix A.3). Our main evaluation and analysis uses 2-shot for higher performance, and 0-shot results can be found in Appendix B.2.
@@ 4.1 Evaluation results
Table 1 summarizes the en-x and x-x MRC performances of the main-list models (see more results in Appendix B.2).
@@ 4.1.1 Cross-lingual performance
Generally, the English MRC performance of most models are high (over 70 out of 100), but the en-x scores ranges (from 45 to 78), showing the performance gap in context retrieval with English and non-English queries. Down into individual models, while GPT-4o shows the highest cross-lingual performance and smallest language gap, several post-trained open LLMs, such as Gemma-2-IT, Qwen-2.5-Instruct and LLaMA-3.1-Instruct, show performance levels and small language gaps comparable to the commercial models.
An interesting observation is that, for LLaMA-3.1 and Gemma-2, the en-en performances remains close after post-training, but en-x greatly improve. This phenomenon is more prominent in smaller (7-9B) than larger models, bring the former a smaller performance gap between English and non-English, which is also observed on Qwen2.5 with growing parameter sizes (See Appendix B.3).
@@ 4.1.2 Comparison with Monolingual performance
In general, the performance gaps between en-en and x-x are much smaller for most models than en-x, and the Qwen models even show higher non-English performance than English. This suggests that non-English language fluency is not the main challenge of xMRC.
Also, for base models and post-trained larger models ($\sim$70B), the x-x performances are always higher than en-x. A possible explanation to this may be the cross-lingual task is less frequent in training, and more difficult because it requires cross-lingual understanding.
However, for post-trained, smaller models (7-9B), the pattern flips, where x-x performances become consistently lower than en-x, suggesting that these models use their English context processing ability to assist non-English retrieval, overcoming the difficulty and low-frequency of the cross-lingual task. Also, since larger LLMs tend to be better at instruction following and understanding, this further highlights that post-training better elicits the cross-lingual context retrieval ability on smaller LLMs.
@@ 4.2.1 Error type ablation
The right part of Table 1 shows error rates of different types (see details in Appendix B.4).
Language and generation errors.
The language error rates of post-trained models (lower part of the table) are significantly lower than base models (upper part of the table), since the former can better follow the cross-lingual task format. Meanwhile, the error rate is low for all the post-trained models, so it cannot be viewed as a bottleneck of xMRC. The generation error rates are minor for most models, regardless of size and post-training, marking it not the bottleneck of xMRC either.